# Localized statistics decoding: A parallel decoding algorithm for quantum low-density parity-check codes

### v1

@@ 1 Introduction
A significant limitation of the BP+OSD decoder is its large runtime overhead. This inefficiency stems primarily from the OSD algorithm’s inversion step, which relies on Gaussian elimination and has cubic worst-case time complexity in the size of the corresponding check matrix. In practice, this is a particularly acute problem, as decoders must be run on large circuit-level decoding graphs that account for errors occurring at any location in the syndrome extraction circuit. This shortcoming constitutes a known barrier to the experimental implementation of quantum LDCP codes, as circuit-level decoding graphs can contain tens of thousands of nodes [12]. Even with specialized hardware, inverting the matrix of a graph of this size cannot realistically be achieved within the decoherence time of a typical qubit [27]. Whilst the BP+OSD decoder is a useful tool for simulations, it is not generally considered a practical method for real-time decoding.
@@ 2.1 The decoding problem
For the decoding of QLDPC codes, the decoder is provided with a matrix $H\in\mathbbm{F}_{2}^{|D|\times|F|}$ called the detector check matrix. This matrix maps circuit fault locations $F$ to so-called detectors $D$, defined as linear combinations of check measurement outcomes that are deterministic in the absence of errors. Specifically, each row of $H$ corresponds to a detector and each column to a fault, and $H_{df}=1$ if fault $f\in\{1,\dots,|F|\}$ flips detector $d\in\{1,\dots,|D|\}$. Such a check matrix can be constructed by tracking the propagation of errors through the syndrome extraction circuit using a stabilizer simulator [30, 31].
We emphasize that, once the detector matrix $H$ is created, the decoding problem can be mapped to the problem of decoding a classical linear code. Given a syndrome $\mathbf{s}\in\mathbbm{F}_{2}^{|D|}$, the decoding problem consists of finding a minimum-weight recovery $\hat{\mathbf{e}}$ such that $\mathbf{s}=H\cdot\hat{\mathbf{e}}$, where the vector $\hat{\mathbf{e}}\in\mathbbm{F}_{2}^{|F|}$ indicates the locations in the circuit where faults have occurred.
The decoding graph is a bipartite graph $\mathcal{G}(H)=(V_{D}\cup V_{F},E)$ with detector nodes $V_{D}$, fault nodes $V_{F}$ and edges $(d,f)\in E\iff H_{df}=1$. $\mathcal{G}(H)$ is also known as the Tanner graph of the check matrix $H$. Since the detector check matrix is analogous to a parity check matrix of a classical linear code, we use the terms detectors and checks synonymously.
@@ 2.2 Localized statistics decoding
This section provides an example-guided outline of the localized statistics decoder. A more formal treatment, including pseudo-code, can be found in Section 4.1.
@@
The localized statistics decoding (LSD) algorithm belongs to the class of reliability-guided inversion decoders, which also contains ordered statistics decoding (OSD) [26, 11, 21]. These solve the decoding problem by computing $\mathbf{\hat{e}}_{[I]}=H_{[I]}^{-1}\cdot\mathbf{s}$. Here, $H_{[I]}$ is an invertible matrix formed by selecting a linearly independent subset of the columns of the check matrix $H$ indexed by the set of column indices $I$. The algorithm is reliability-guided in that it uses prior knowledge of the error distribution to strategically select $I$ so that the solution $\mathbf{\hat{e}}_{[I]}$ spans faults that have the highest error probability. The reliabilities can be derived, for example, from the device’s physical error model [32, 33, 34, 35, 16, 36] or the soft information output of a pre-decoder such as BP [37].
@@
Figure 1 shows an example of error factorization in the Tanner graph of a $5\times 10$ surface code. The support of a fault vector $\mathbf{e}$ is illustrated by the circular nodes marked with an $X$ and the corresponding syndrome is depicted by the square nodes filled in red. In this example, it is clear that $\mathbf{e}$ can be split into two connected components, $\mathbf{e}_{[C_{1}]}$ and $\mathbf{e}_{[C_{2}]}$, that occupy separate regions of the decoding graph. We refer to each of the connected components induced by an error on the decoding graph as clusters. With a slight misuse of notation, we refer to clusters $C_{i}$ and their associated incidence matrices $H_{[C_{i}]}$ interchangeably. This identification is natural as clusters are uniquely identified by their fault nodes, or equivalently, by column indices of $H$: for a set of fault nodes $C\subseteq V_{F}$, we consider all of the detector nodes in $V_{D}$ adjacent to at least one node in $C$ to be part of the cluster.
@@
$\mathbf{\hat{e}}_{[C_{1}\cup C_{2}\cup C_{\bot}]}=\left(H_{[C_{1}]}^{-1}\mathbf{s}_{[C_{1}]},H_{[C_{2}]}^{-1}\mathbf{s}_{[C_{2}]},0\right)\rm,$ (1)
@@
$\mathbf{\hat{e}}_{[C_{1}\cup\dots\cup C_{\nu}\cup C_{\bot}]}=\left(H_{[C_{1}]}^{-1}\mathbf{s}_{[C_{1}]},\dots,H_{[C_{\nu}]}^{-1}\mathbf{s}_{[C_{\nu}]},0\right)\rm.$ (2)
The number of clusters, $\nu$, will depend upon $H$, the physical error rate, and the Hamming weight of $\mathbf{s}$. If a factorization can be found, matrix inversion is efficient: first, the $\nu$ clusters can be solved in parallel; second, the parallel worst-case time complexity of the algorithm depends on the maximum size of a cluster $\kappa=\text{max}_{i}\left(|C_{i}|\right)$, where $|C_{i}|$ is the number of fault nodes in $C_{i}$. The worst-case scaling $O(\kappa^{3})$ contrasts with the $O(n^{3})$ OSD post-processing scaling, where $n=|V_{F}|$ is the size of the matrix $H$. To enable parallel execution, we have devised a routine that we call on-the-fly elimination to efficiently merge clusters and compute a matrix factorization, as detailed in Section 4.2.
@@
The LSD decoder uses a weighted, reliability-based growth strategy to factorize the decoding graph. The algorithm begins by creating a cluster $H_{[C_{i}]}$ for each flipped detector node, i.e., a separate cluster is generated for every nonzero bit in the syndrome vector $\mathbf{s}$. At every growth step, each cluster $H_{[C_{i}]}$ is grown by one column by adding the fault node from its neighborhood that has the highest probability of being in error according to the input reliability information. This weighted growth strategy is particularly important for controlling the cluster size: limiting growth to a single fault node per time step ensures that an efficient factorization can be found, even for QLDPC codes with high degrees of expansion in their decoding graphs.
@@
The process of weighted-cluster growth is illustrated in Fig. 2 for the surface code. Here, two clusters are created. These are grown according to the reliability ordering of the neighboring fault nodes. In Fig. 2(c), the two clusters merge, yielding a combined valid cluster. The combined cluster is not optimal as its associated decoding matrix has $5$ columns, whereas the local solution has Hamming weight $3$, indicating that the optimal cluster would have $3$ columns. Nonetheless, computing a solution using the cluster matrix with $5$ columns is still preferable to computing a solution using the full $41$-column decoding matrix – this highlights the possible computational gain of LSD.
@@
As a timely example, we focus specifically on the cluster size statistics of the circuit-noise decoding graph of the $\llbracket 144,12,12\rrbracket$ bivariate bicycle code [39] that was recently investigated in Ref. [12]. Fig. 3(a) shows the distribution of the maximum sizes of clusters identified by BP+LSD over $10^{5}$ decoding samples, see Section 4.3 for details. The figure illustrates that for low enough noise rates, the largest clusters found by LSD are small and close to the optimal sizes of clusters induced by the original error, even if only a relatively small number (30) of BP iterations is used to compute the soft information input to LSD. It is worth emphasizing that large clusters are typically formed by merging two, or more, smaller clusters identified and processed at previous iterations of the algorithm. Owing to our on-the-fly technique that processes the linear system corresponding to each of these clusters (cf. Section 4.2), the maximum cluster size only represents a loose upper bound on the complexity of the LSD algorithm.
Fig. 3(b) shows the distributions of the cluster count per error, $\nu$, against the physical error rate $p$. The number of clusters $\nu$ corresponds to the number of terms in the factorization of the decoding problem and thus indicates the degree to which the decoding can be parallelized, as disjoint factors can be solved concurrently. At practically relevant error rates below the (pseudo) threshold, e.g., $p\leq 0.1\%$, we observe on average 10 independent clusters. This implies that the LSD algorithm benefits from parallel resources throughout its execution.
@@
Figure 3: Cluster size statistics of the $\llbracket 144,12,12\rrbracket$ bivariate bicycle code of Ref. [12] under circuit-level noise with strength $p$. Markers show the mean of the distribution while shapes are violin plots of the distribution obtained from $10^{5}$ samples. Yellow distributions show statistics for the optimal factorization while the blue distributions show statistics for the factorization returned by BP+LSD. We show in panel (a) the distribution of the maximum cluster size $\kappa$ and in (b) the distribution of the cluster count, $\nu$, for each decoding sample. Markers and distributions are slightly offset from the actual error rate to increase readability.
@@
[…]
Figure 4: Comparison of various decoders guided by belief propagation for decoding rotated surface codes of distance $d$ subject to circuit-level depolarizing noise parameterized by a single parameter, called the physical error rate $p$, see Section 4.3 for details. We use Stim to perform a surface_code:rotated_memory_z experiment for $d$ syndrome extraction cycles with single and two-qubit error probabilities $p$. (a) The performance of BP+OSD-0 that matches the performance of the proposed decoder. (b) The performance of a BeliefFind decoder that shares a cluster growth strategy with the proposed decoder. (c) Performance of the proposed BP+LSD decoder. The shading indicates hypotheses whose likelihoods are within a factor of 1000 of the maximum likelihood estimate, similar to a confidence interval.
@@ 2.3 Numerical results
Our main numerical finding is that BP+LSD can decode QLDPC codes with performance on par with BP+OSD. We include the results of extensive simulations in which BP+LSD is used to decode a circuit-level depolarizing noise model for surface codes, hypergraph product codes [41], and bivariate bicycle codes [12, 39].
@@
Figure 5: Below threshold logical error rate $p_{L}$ of a family of $\llbracket 25s^{s},s^{2}\rrbracket$ constant-rate hypergraph product codes decoded with of the BP+LSD decoder. We simulate $N_{c}=12$ rounds of syndrome extraction cycles under circuit-level noise with physical error rate $p$ and apply a $(3,1)-$overlapping window technique to enable fast and accurate single-shot decoding, see Section 4.3 for details. The shading indicates hypotheses whose likelihoods are within a factor of 1000 of the maximum likelihood estimate, similar to a confidence interval. Dashed lines are a fit to the numerically observed error rates.
@@ 3 Discussion
When considering large QLDPC codes, current state-of-the-art decoders such as BP+OSD hit fundamental limitations due to the size of the resulting decoding graphs. This limitation constitutes a roadblock in the realization of protocols based on QLDPC codes. In this work, we address this challenge through the introduction of the LSD decoder as a parallel algorithm whose runtime depends predominantly on the physical error rate of the system. Our algorithm uses a reliability-based growth procedure to construct clusters on the decoding graph in a parallel fashion. Using a novel routine that computes the PLU decomposition of the clusters’ sub-matrices on-the-fly, we can merge clusters efficiently and compute local decoding solutions in a parallel fashion. Our main numerical findings are that the proposed decoder performs on par with current state-of-the-art decoding methods in terms of logical decoding performance.
@@ 4.1 LSD algorithm
$\beta(C)=\{d\in V_{D}^{C}\mid\Gamma(d)\not\subseteq V_{F}^{D}\}$ (3)
@@
A sequential version of the algorithm is outlined below and detailed in the pseudo-code in Algorithm 1. A parallel version of the LSD algorithm is presented in Appendix B.
@@
    Every cluster is grown by a single node $v_{j}$ drawn from the list of candidate nodes $\Lambda(C_{i})$. For the first growth step after cluster initialization, we define $\Lambda(C_{i})=\Gamma(\{s_{i}\})$ – see Eq. 4. The chosen growth node $v_{j}\in\Lambda(C_{i})$ in each step is the fault node with the highest probability of being in error. That is, $v_{j}$ has the lowest value among the LLRs for the candidate fault nodes $\lambda_{j}<\lambda_{j+1}<\dots<\lambda_{\ell},\ \ell=|\Lambda(C)|$. Hence, the growth step involves adding one new column to the cluster matrix $H_{[C_{i}]}$.
@@
    The Gaussian elimination row operations performed on previous columns are performed on the new column together with the row operations needed to eliminate the newly added columns of $H_{[C_{i}]}$. In addition, every row operation applied to $H_{[C_{i}]}$ is also applied to the local syndrome $\mathbf{s}_{[C_{i}]}$. This allows the algorithm to efficiently track when the cluster becomes valid. Explicitly, the cluster is valid when the syndrome becomes linearly dependent on the cluster decoding matrix i.e., when $\mathbf{s}_{[C_{i}]}\in\text{{image}}(H_{[C_{i}]})$. In addition to cluster validation, the Gaussian elimination at each step enables an on-the-fly computation of the PLU factorization of the local cluster. We refer the reader to Section 4.2 for an outline of our on-the-fly elimination method.

### v2

# Localized statistics decoding for quantum low-density parity-check codes

@@ 1 Introduction
A significant limitation of the BP+OSD decoder is its large runtime overhead. This inefficiency stems primarily from the OSD algorithm’s inversion step, which relies on Gaussian elimination and has cubic worst-case time complexity in the size of the corresponding check matrix. In practice, this is a particularly acute problem, as decoders must be run on large circuit-level decoding graphs that account for errors occurring at any location in the syndrome extraction circuit. This shortcoming constitutes a known barrier to the experimental implementation of efficient quantum LDCP codes, as circuit-level decoding graphs can contain tens of thousands of nodes [12]. Even with specialized hardware, inverting the matrix of a graph of this size cannot realistically be achieved within the decoherence time of a typical qubit [27]. Whilst the BP+OSD decoder is a useful tool for simulations, it is not generally considered a practical method for real-time decoding.
@@ 2.1 The decoding problem
For the decoding of QLDPC codes, the decoder is provided with a matrix $H\in\mathbbm{F}_{2}^{|D|\times|F|}$ called the detector check matrix. This matrix maps circuit fault locations $F$ to so-called detectors $D$, defined as linear combinations of check measurement outcomes that are deterministic in the absence of errors. Specifically, each row of $H$ corresponds to a detector and each column to a fault, and $H_{df}=1$ if fault $f\in\{1,\dots,|F|\}$ flips detector $d\in\{1,\dots,|D|\}$. Such a check matrix can be constructed by tracking the propagation of errors through the syndrome extraction circuit using a stabilizer simulator [30, 31, 32].
We emphasize that, once the detector matrix $H$ is created, the minium-weight decoding problem can be mapped to the problem of decoding a classical linear code: Given a syndrome $\mathbf{s}\in\mathbbm{F}_{2}^{|D|}$, the decoding problem consists of finding a minimum-weight recovery $\hat{\mathbf{e}}$ such that $\mathbf{s}=H\cdot\hat{\mathbf{e}}$, where the vector $\hat{\mathbf{e}}\in\mathbbm{F}_{2}^{|F|}$ indicates the locations in the circuit where faults have occurred.
The decoding graph is a bipartite graph $\mathcal{G}(H)=(V_{D}\cup V_{F},E)$ with detector nodes $V_{D}$, fault nodes $V_{F}$ and edges $(d,f)\in E\iff H_{df}=1$. $\mathcal{G}(H)$ is also known as the Tanner graph of the check matrix $H$. Since the detector check matrix is analogous to a parity check matrix of a classical linear code, we use the terms detectors and checks synonymously. Note that we implement a minimum-weight decoding strategy where the goal is to find the lowest-weight error compatible with the syndrome. This is distinct from maximum-likelihood decoding, where the goal is to determine the highest probability logical coset.
@@ 2.2 Localized statistics decoding
This section provides an example-guided outline of the localized statistics decoder. A more formal treatment, including pseudo-code, can be found in Section˜4.1.
@@
The localized statistics decoding (LSD) algorithm belongs to the class of reliability-guided inversion decoders, which also contains ordered statistics decoding (OSD) [26, 11, 21]. OSD can solve the decoding problem by computing $\mathbf{\hat{e}}_{[I]}=H_{[I]}^{-1}\cdot\mathbf{s}$. Here, $H_{[I]}$ is an invertible matrix formed by selecting a linearly independent subset of the columns of the check matrix $H$ indexed by the set of column indices $I$. The algorithm is reliability-guided in that it uses prior knowledge of the error distribution to strategically select $I$ so that the solution $\mathbf{\hat{e}}_{[I]}$ spans faults that have the highest error probability. The reliabilities can be derived, for example, from the device’s physical error model [33, 34, 35, 36, 16, 37] or the soft information output of a pre-decoder such as BP [38].
@@
Figure˜1 shows an example of error factorization in the Tanner graph of a $5\times 10$ surface code. The support of a fault vector $\mathbf{e}$ is illustrated by the circular nodes marked with an $X$ and the corresponding syndrome is depicted by the square nodes filled in red. In this example, it is clear that $\mathbf{e}$ can be split into two connected components, $\mathbf{e}_{[C_{1}]}$ and $\mathbf{e}_{[C_{2}]}$, that occupy separate regions of the decoding graph. We refer to each of the connected components induced by an error on the decoding graph as clusters. With a slight misuse of notation, we refer to clusters $C_{i}$ and their associated incidence matrices $H_{[C_{i}]}$ interchangeably and use $|C_{i}|$ to denote the number of fault nodes (columns) in the cluster (its incidence matrix, respectively). This identification is natural as clusters are uniquely identified by their fault nodes, or equivalently, by column indices of $H$: for a set of fault nodes $C\subseteq V_{F}$, we consider all of the detector nodes in $V_{D}$ adjacent to at least one node in $C$ to be part of the cluster.
@@
$\mathbf{\hat{e}}_{[C_{1}\cup C_{2}\cup C_{\bot}]}=\left(H_{[C_{1}]}^{-1}\mathbf{s}_{[C_{1}]},H_{[C_{2}]}^{-1}\mathbf{s}_{[C_{2}]},0\right),$ (1)
@@
$\mathbf{\hat{e}}_{[C_{1}\cup\dots\cup C_{\nu}\cup C_{\bot}]}=\left(H_{[C_{1}]}^{-1}\mathbf{s}_{[C_{1}]},\dots,H_{[C_{\nu}]}^{-1}\mathbf{s}_{[C_{\nu}]},0\right).$ (2)
The number of clusters, $\nu$, will depend upon $H$, the physical error rate, and the Hamming weight of $\mathbf{s}$. If a factorization can be found, matrix inversion is efficient: first, the $\nu$ clusters can be solved in parallel; second, the parallel worst-case time complexity of the algorithm depends on the maximum size of a cluster $\kappa=\text{max}_{i}\left(|C_{i}|\right)$, where $|C_{i}|$ is the number of fault nodes in $C_{i}$. The worst-case scaling $O(\kappa^{3})$ contrasts with the $O(n^{3})$ OSD post-processing scaling, where $n=|V_{F}|$ is the size of the matrix $H$. To enable parallel execution, we have devised a routine that we call on-the-fly elimination to efficiently merge clusters and compute a matrix factorization, as detailed in Section˜4.2.
@@
The LSD decoder uses a weighted, reliability-based growth strategy to factorize the decoding graph. The algorithm begins by creating a cluster $H_{[C_{i}]}$ for each flipped detector node, i.e., a separate cluster is generated for every nonzero bit in the syndrome vector $\mathbf{s}$. At every growth step, each cluster $H_{[C_{i}]}$ is grown by one column by adding the fault node from its neighborhood with the highest probability of being in error according to the input reliability information. This weighted growth strategy is crucial for controlling the cluster size: limiting growth to a single fault node per time step increases the likelihood that an efficient factorization is found, especially for QLDPC codes with high degrees of expansion in their decoding graphs.
@@
The process of weighted-cluster growth is conceptually similar to “belief hypergraph union-find” [39] and is illustrated in Fig. 2 for the surface code. Here, two clusters are created. These are grown according to the reliability ordering of the neighboring fault nodes. In Fig. 2(c), the two clusters merge, yielding a combined valid cluster. The combined cluster is not optimal as its associated decoding matrix has $5$ columns, whereas the local solution has Hamming weight $3$, indicating that the optimal cluster would have $3$ columns. Nonetheless, computing a solution using the cluster matrix with $5$ columns is still preferable to computing a solution using the full $41$-column decoding matrix – this highlights the possible computational gain of LSD.
@@
As a timely example, we focus specifically on the cluster size statistics of the circuit-noise decoding graph of the $\llbracket 144,12,12\rrbracket$ bivariate bicycle code [41] that was recently investigated in Ref. [12]. Fig. 3(a) shows the distribution of the maximum sizes of clusters identified by BP+LSD over $10^{5}$ decoding samples, see Section˜4.3 for details. The figure illustrates that for low enough noise rates, the largest clusters found by LSD are small and close to the optimal sizes of clusters induced by the original error, even if only a relatively small number (30) of BP iterations is used to compute the soft information input to LSD. It is worth emphasizing that large clusters are typically formed by merging two, or more, smaller clusters identified and processed at previous iterations of the algorithm. Owing to our on-the-fly technique that processes the linear system corresponding to each of these clusters (cf. Section˜4.2), the maximum cluster size only represents a loose upper bound on the complexity of the LSD algorithm.
Fig. 3(b) shows the distributions of the cluster count per shot, $\nu$ – that is, per shot, where the LSD data is post-selected on shots where BP does not converge – against the physical error rate $p$. The number of clusters $\nu$ corresponds to the number of terms in the factorization of the decoding problem and thus indicates the degree to which the decoding can be parallelized, as disjoint factors can be solved concurrently. At practically relevant error rates below the (pseudo) threshold, e.g., $p\leq 0.1\%$, we observe on average 10 independent clusters. This implies that the LSD algorithm benefits from parallel resources throughout its execution.
@@
Figure 3: Cluster size statistics of the $\llbracket 144,12,12\rrbracket$ bivariate bicycle code of Ref. [12] under circuit-level noise with strength pp. Markers show the mean of the distribution while shapes are violin plots of the distribution obtained from $10^{5}$ samples. Yellow distributions show statistics for the optimal factorization while the blue distributions show statistics for the factorization returned by BP+LSD. We show in panel (a) the distribution of the maximum cluster size $\kappa$ and in (b) the distribution of the cluster count, $\nu$, for each decoding sample. Markers and distributions are slightly offset from the actual error rate to increase readability.
@@
Higher-order reprocessing in OSD is a systematic approach designed to increase the decoder’s accuracy. The zero-order solution $\mathbf{\hat{e}}_{[I]}=H_{[I]}^{-1}\cdot\mathbf{s}$ of the decoder cannot be made lower weight if the set of column indices $I$ specifying the invertible submatrix $H_{[I]}$ matches the $|I|$ most likely fault locations identified from the soft information vector $\mathbf{\lambda}$. However, if there are linear dependencies within the columns formed by the $|I|$ most likely fault locations, the solution $\hat{\mathbf{e}}$ may not be optimal. In those cases, some fault locations in $\overline{I}$ (the complement of $I$) might have higher error probabilities. To find the optimal solution, one can systematically search all valid fault configurations in $\overline{I}$ that potentially provide a more likely estimate $\hat{\mathbf{e}}^{\prime}$. This search space, however, is exponentially large in $|\overline{I}|$. Thus, in practice, only configurations with a Hamming weight up to $w$ are considered, known as orderw-w reprocessing. See Refs. [25, 26, 11, 21] for a more technical discussion.
@@
Figure 4: Comparison of various decoders guided by belief propagation for decoding rotated surface codes of distance $d$ subject to circuit-level depolarizing noise parameterized by a single parameter, called the physical error rate $p$, see Section˜4.3 for details. We use Stim to perform a surface_code:rotated_memory_z experiment for $d$ syndrome extraction cycles with single and two-qubit error probabilities pp. (a) The performance of BP+OSD-0 that matches the performance of the proposed decoder. (b) The performance of a BeliefFind decoder that shares a cluster growth strategy with the proposed decoder. (c) Performance of the proposed BP+LSD decoder. The shading indicates hypotheses whose likelihoods are within a factor of 1000 of the maximum likelihood estimate, similar to a confidence interval.
@@ 2.3 Numerical results
Our main numerical finding is that BP+LSD can decode QLDPC codes with performance on par with BP+OSD. We include the results of extensive simulations in which BP+LSD is used to decode a circuit-level depolarizing noise model for surface codes, hypergraph product (HGP) codes [43], and bivariate bicycle codes [12, 41].
@@
Figure 5: Below threshold logical error rate $p_{L}$ of a family of $\llbracket 25s^{2},s^{2}\rrbracket$ constant-rate hypergraph product codes decoded with of the BP+LSD decoder. We simulate $N_{c}=12$ rounds of syndrome extraction cycles under circuit-level noise with physical error rate $p$ and apply a $(3,1)$-overlapping window technique to enable fast and accurate single-shot decoding, see Section˜4.3 for details. The shading indicates hypotheses whose likelihoods are within a factor of 1000 of the maximum likelihood estimate, similar to a confidence interval. Dashed lines are an exponential fit with a linear exponent to the numerically observed error rates.
@@
Runtime statistics.
To estimate the time overhead of the proposed decoder in numerical simulation scenarios and to compare it with a state-of-the-art implementation of BP+OSD, we present preliminary timing results for our prototypical open-source implementation of LSD in Appendix C.4. We note that for a more complete assessment of performance, it will be necessary to benchmark a fully parallel implementation of the algorithm, designed for specialized hardware such as GPUs or FPGAs. We leave this as a topic for future work.
@@ 3 Discussion
When considering large QLDPC codes, current state-of-the-art decoders such as BP+OSD hit fundamental limitations due to the size of the resulting decoding graphs. This limitation constitutes a severe bottleneck in the realization of protocols based on QLDPC codes. In this work, we address this challenge through the introduction of the LSD decoder as a parallel algorithm whose runtime depends predominantly on the physical error rate of the system. Our algorithm uses a reliability-based growth procedure to construct clusters on the decoding graph in a parallel fashion. Using a novel routine that computes the PLU decomposition [46] of the clusters’ sub-matrices on-the-fly, we can merge clusters efficiently and compute local decoding solutions in a parallel fashion. Our main numerical findings are that the proposed decoder performs on par with current state-of-the-art decoding methods in terms of logical decoding performance.
@@ 4.1 LSD algorithm
$\beta(C)=\{d\in V_{D}^{C}\mid\Gamma(d)\not\subseteq V_{F}^{C}\}$ (3)
@@
A sequential version of the algorithm is outlined below and detailed in the pseudo-code in Algorithm˜1. A parallel version of the LSD algorithm is presented in Appendix B.
@@
    Every cluster is grown by a single node $v_{j}$ drawn from the list of candidate nodes $\Lambda(C_{i})$. For the first growth step after cluster initialization, we define $\Lambda(C_{i})=\Gamma(\{s_{i}\})$ – see Eq.˜4. The chosen growth node $v_{j}\in\Lambda(C_{i})$ in each step is the fault node with the highest probability of being in error. That is, $v_{j}$ has the lowest value among the LLRs for the candidate fault nodes $\lambda_{j}<\lambda_{j+1}<\dots<\lambda_{\ell},\ \ell=|\Lambda(C)|$. Hence, the growth step involves adding one new column to the cluster matrix $H_{[C_{i}]}$.
@@
    The Gaussian elimination row operations performed on previous columns are performed on the new column together with the row operations needed to eliminate the newly added columns of $H_{[C_{i}]}$. In addition, every row operation applied to $H_{[C_{i}]}$ is also applied to the local syndrome $\mathbf{s}_{[C_{i}]}$. This allows the algorithm to efficiently track when the cluster becomes valid. Explicitly, the cluster is valid when the syndrome becomes linearly dependent on the cluster decoding matrix i.e., when $\mathbf{s}_{[C_{i}]}\in\text{{image}}(H_{[C_{i}]})$. In addition to cluster validation, the Gaussian elimination at each step enables an on-the-fly computation of the PLU factorization of the local cluster. We refer the reader to Section˜4.2 for an outline of our on-the-fly elimination method.